OpenStudy (anonymous):

intergal (2x-3)5dx can any one help me to solve this problem????

OpenStudy (anonymous):

is it \[\int(2x-3)5dx\]?

OpenStudy (anonymous):

is so multiply out and start with \[\int(10x-15)dx\]

OpenStudy (anonymous):

5x^2-15x+c

OpenStudy (beginnersmind):

I'm guessing it's (2x-3)^5 dx If so the substitution u=2x-3 should work.

OpenStudy (anonymous):

i already did that and i also divided 2 in both sides

OpenStudy (anonymous):

i came up with 1/2du=dx

OpenStudy (beginnersmind):

Just to clarify, the integral was: \[\int\limits(2x-3)^{5}dx\] You substituted u = (2x-3) and dx=1/2 du (which is correct). So you should have \[\int\limits \frac{1}{2} u^{5}du\]

OpenStudy (anonymous):

@beginnersmind : you are right if (2x-5)^5 chose u=2x-5 du=2dx 1/2 du=dx integral 1/2 u du 1/4 u^2+c 1/4(2x-5)^2+c

OpenStudy (beginnersmind):

edit: du=1/2 dx that is.

OpenStudy (anonymous):

is it true buddy?

OpenStudy (anonymous):

ummm not really because that's not the answer

OpenStudy (anonymous):

yes of course silly me

OpenStudy (beginnersmind):

don't worry about the answer yet. Rick made some errors in calculation. Is the last line in my previous post the same what you got so far?

OpenStudy (anonymous):

but you should have \[\frac{1}{2}\int u^5du=\frac{1}{12}u^6\]

OpenStudy (anonymous):

raise the power by one and divide by it power rule backwards

OpenStudy (anonymous):

I think we integrate it well.. so what is the answer?

OpenStudy (anonymous):

beginnersmind yes so far your steps are right ? keep on going

OpenStudy (anonymous):

f \[\frac{1}{12}u^6\] substitute back get \[\frac{1}{12}(2x-5)^6\]

OpenStudy (anonymous):

Rizki the answer is (2x-3)^6/12+c

OpenStudy (beginnersmind):

You integrate with respect to u. As satellite said, it's the power rule backwards. Then substitute back (2x-3) for u. Don't forget the integration constant either ;)

OpenStudy (anonymous):

i'm sorry, i took u^2 not u^5. Yes, you are right.

OpenStudy (anonymous):

beginnersmind can u finish it up plzz

OpenStudy (beginnersmind):

satellite basically did. Just follow his steps, fix the typos and don't forget to add the constant.

OpenStudy (anonymous):

yeah always some typos for sure

OpenStudy (anonymous):

[1/2\int\limits U^3 dx where that come from ?

OpenStudy (beginnersmind):

\[\frac{1}{2}\int\limits\ u^{3} du \] anywhere. Did you mean \[\frac{1}{2}\int\limits\ u^{5} du \] ?

OpenStudy (anonymous):

oooh ok i gat it now i was making silly mistake

OpenStudy (anonymous):

thank you soo much guys i really appreicate the help

OpenStudy (beginnersmind):

No problem :)